# Semitoric Families on Pentagon Spaces

### v1

@@ 2. Polygon spaces
Let $\widetilde{M}:=S^{2}_{r_{1}}\times\cdots\times S^{2}_{r_{n}}$ be the product of symplectic $2$-spheres of radius $r_{1},\ldots,r_{n}$ and let $\mu:\widetilde{M}\to{\mathbb{R}}^{3}$ be the moment map of the diagonal Hamiltonian $SO(3)$-action on $\widetilde{M}$. Under this setting, $\mathcal{P}(r_{1},\ldots,r_{n})=\widetilde{M}\sslash_{0}SO(3)=\mu^{-1}(0)/SO(3)$. Notice that there is a Hamiltonian $S^{1}$-action on $\widetilde{M}$ with moment map $\widetilde{\ell_{12}}(\rho)=|\rho_{1}+\rho_{2}|$, and this $S^{1}$-action commutes with the $SO(3)$-action on $\widetilde{M}$. Hence, we can identify $\ell_{12}^{-1}(c)/S^{1}$ with $F^{-1}(c,0)/(S^{1}\times SO(3))$, where $F=(\widetilde{\ell_{12}},\mu):\widetilde{M}\to{\mathbb{R}}\times{\mathbb{R}}^{3}$. It follows that we can do symplectic reductions by stages: first by the $S^{1}$-action and then by the $SO(3)$-action. In other words, $\ell_{12}^{-1}(c)/S^{1}=(\widetilde{M}\sslash_{c}S^{1})\sslash_{0}SO(3)$.
@@
Let $I,J$ be nonempty subsets of $\{1,\ldots,n\}$. If one of the following holds
@@
We next establish a general criterion characterizing the singular points of $H_{t}$ that will be used in our semitoric families.
@@
Fix $p\in\mathcal{P}(r_{1},\ldots,r_{n})$ and choose a representative $\rho\in\widetilde{\mathcal{P}}(r_{1},\ldots,r_{n})$ of $p$ with $\rho_{1}=(r_{1},0,0)$. For any $t\in[0,1]$, $p$ is a singular point of $H_{t}:=t\ell_{n-2,n-1}^{2}+(1-t)\ell^{2}_{n-1,n}$ if and only if there exists an $a\in{\mathbb{R}}$ such that
  • $(a,0,0)\times\rho_{i}=0$ for $i\neq n-2,n-1,n$,
  • $(a,0,0)\times\rho_{n-2}=t\rho_{n-1}\times\rho_{n-2}$, and
@@
Let $r_{1},\ldots,r_{5}$ be positive numbers such that
  • $r_{4}<r_{5}<r_{3}$
  • $|r_{1}+r_{2}-r_{3}|<r_{5}-r_{4}$
  • $|r_{3}-r_{4}-r_{5}|<|r_{1}-r_{2}|<r_{3}+r_{4}-r_{5}$
Then the $S^{1}$-action on $M:=\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ with moment map $J=\ell_{12}$ is semi-free such that $M^{S^{1}}$ has three connected components: two fixed surfaces $J^{-1}(|r_{1}-r_{2}|),J^{-1}(r_{1}+r_{2})$ and an isolated fixed point $p\in M$ with $J(p)=r_{3}+r_{4}-r_{5}$. Moreover, $p$ is also a singular point of $H_{t}:=t\ell_{34}^{2}+(1-t)\ell_{45}^{2}$ for any $t\in[0,1]$.
@@
By Lemma 2.3 and [KM96, Proposition 3.11] , $(\ell_{12},\ell_{34})$ (or $(\ell_{12},\ell_{45})$) is the moment map of an effective Hamiltonian $T^{2}$-action on $M$. It is straightforward to check that the Delzant polygons corresponding to this toric action are given by Figure 1. The claims about fixed components follow directly from Delzant’s classification [Del88] and the local normal form theorem [GS84]. Let $p\in M$ be the preimage of $(r_{3}+r_{4}-r_{5},r_{3}+r_{4})$ and let $\rho\in\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})$ be a representative of $p$. Since $\ell_{34}(p)=r_{3}+r_{4}$, we know that $\rho_{3}\times\rho_{4}=0$. Since $\rho_{1}+\rho_{2}=-\rho_{3}-\rho_{4}-\rho_{5}$ and $\ell_{12}(p)=r_{3}+r_{4}-r_{5}$, we know that $\rho_{4}\times\rho_{5}=0$ as well. Hence, by Lemma 2.4, $p$ is a singular point of $H_{t}$ for all $t\in[0,1]$.
$(J_{\min},r_{3}-r_{4})$$(J_{\max},r_{3}-r_{4})$$(J_{\max},r_{3}+r_{4})$$(j,r_{3}+r_{4})$$(J_{\min},r_{5}+J_{\min})$
(a) moment image of $(\ell_{12},\ell_{34})$
$(J_{\min},r_{4}+r_{5})$$(J_{\max},r_{4}+r_{5})$$(J_{\max},r_{5}-r_{4})$$(j,r_{5}-r_{4})$$(J_{\min},r_{3}-J_{\min})$
(b) moment image of $(\ell_{12},\ell_{45})$
Figure 1. In both figures, $J_{\min}=|r_{1}-r_{2}|,J_{\max}=r_{1}+r_{2},j=r_{3}+r_{4}-r_{5}$.
@@ 3. Semitoric Systems
In this section, we recall the definition of semitoric systems and the classification of non-degenerate singularities. The non-degeneracy of singularities can be defined in any dimension for any integrable system. Since our main interest lies in dimension 4, we will follow Section 2.3 in [LFP23] to give the equivalent definition using Hessian matrices. For the definition in any dimension using Cartan subalgebras, we refer the readers to [BF04, Definition 1.23].
@@ 3.1. Singularities of $4$-dimensional integrable systems
  • hyperbolic-hyperbolic if $\lambda_{1},\lambda_{2}>0$;
@@
Let $p\in M$ be a rank one singular point. In this case, $dJ_{p}$ and $dH_{p}$ are linearly dependent, so the Hamiltonian vector fields $X_{J}(p),X_{H}(p)$ are linearly dependent as well because of the non-degeneracy of $\omega$. Let $L$ be the real span of $X_{J}(p),X_{H}(p)$ and $L^{\omega}$ be the symplectic orthogonal of $L$. Then $A_{\nu,\mu}:=\Omega_{p}^{-1}(\nu\operatorname{Hess}(J)_{p}+\mu\operatorname{Hess}(H)_{p})$ descends to an operator, by a slight abuse of notation, still called $A_{\nu,\mu}$ on the two-dimensional quotient $L^{\omega}/L$, and the eigenvalues of $A_{\nu,\mu}$ are of the form $\pm\lambda$ for $\lambda\in{\mathbb{C}}$ (again, see Proposition 1.2 in [BF04]).
@@
    $J$ is the moment map for an effective Hamiltonian $S^{1}$-action, and
@@ 3.2. Local normal forms of singularities of a semitoric system
In this section, we recall the local theory of singularities of a semitoric system. The best-known local theory is proposed by Eliasson [Eli84]: it is a symplectic Morse lemma for any integrable system. However, since we are studying semitoric systems only, the local normal form given by the Hamiltonian $S^{1}$-action comes in handy, so we will use the following lemma. We refer the interested readers to Section 2.2 in [LFP23] for a discussion of Eliasson’s normal form and the references therein.
@@ 3.2.1. Rank zero points
[LFP23, Proposition 7.5] Let $J=\epsilon(\frac{1}{2}|z_{1}|^{2}-\frac{1}{2}|z_{2}|^{2})$ with $\epsilon\in\{-1,1\}$. Then $H\in\mathcal{C}^{\infty}({\mathbb{C}}^{2},{\mathbb{R}})$ is a Hamiltonian such that $\{J,H\}=0$, $H(0)=0$ and $dH(0)=0$ if and only if there exist $\mu_{1},\mu_{2},\mu_{3},\psi\in{\mathbb{R}}$ and $R_{3}\in\mathcal{C}^{\infty}({\mathbb{C}}^{2},{\mathbb{R}})$ with $R_{3}(z_{1},z_{2})=O(3)$ and $\{J,R_{3}\}=0$ such that
@@
  • of elliptic-elliptic type if $|\mu_{2}+\mu_{3}|>|\mu_{1}|$;
@@
For a point $p\in M$ fixed by the $S^{1}$-action, the isotropy weights are the weight vectors of the isotropy representation of $S^{1}$ on $T_{p}M$. In dimension $4$, there are two (counted with multiplicity) isotropy weights at each fixed point. The above lemma discusses the case where the isotropy weights are $\pm 1$. In fact, this is the only case where $p$ is possibly a focus-focus singularity of the semitoric system. Another case of interest is where $p$ lies on a surface fixed by the $S^{1}$-action. Since the action is effective, the isotropy weights must be $0,1$ or 0,10,-1. We prove the following lemma assuming that $p$ has isotropy weights $0,1$. Notice that we could have used the complex coordinates $z_{1},z_{2}$ and put $J$ as $|z_{1}|^{2}$, but this will put $H$ in a complicated form. Since the non-degeneracy and the type of singularity only depend on the second-order terms in the Taylor expansion, we use the real coordinates to simplify the statement and the proof.
@@
In this case, we have the following matrices
$\Omega=\begin{pmatrix}\mathbf{J}&\mathbf{O}\\ \mathbf{O}&\mathbf{J}\end{pmatrix},\operatorname{Hess}(J)=\begin{pmatrix}2\mathbf{I}&\mathbf{O}\\ \mathbf{O}&\mathbf{O}\end{pmatrix},\operatorname{Hess}(H)=\begin{pmatrix}2a_{11}\mathbf{I}&\mathbf{O}\\ \mathbf{O}&\mathbf{D}\end{pmatrix}$,
where $\mathbf{J}=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$, $\mathbf{D}=\begin{pmatrix}2a_{22}&b_{22}\\ b_{22}&2c_{22}\end{pmatrix}$, $\mathbf{O}$ is the zero matrix and $\mathbf{I}$ is the identity matrix. Fix a number $\lambda\neq-a_{11}$ and consider $A_{\lambda,1}=\Omega^{-1}(\lambda\operatorname{Hess}(J)+\operatorname{Hess}(H))$. The characteristic polynomial of $A_{\lambda,1}$ is $x^{2}+(2\lambda+2a_{11})^{2}][x^{2}+4a_{22}c_{22}-{b_{22}}^{2}]$. Hence, $\chi_{\lambda,1}(X)=(X+(2\lambda+2a_{11})^{2})(X+4a_{22}c_{22}-{b_{22}}^{2})$. The two roots are two real numbers $\lambda_{1}=-(2\lambda+2a_{11})^{2}$ and $\lambda_{2}={b_{22}}^{2}-4a_{22}c_{22}$. Since $\lambda\neq-a_{11}$, we have that $\lambda_{1}<0$. By Definition 3.1, $(0,0)$ is elliptic-elliptic if and only if $\lambda_{2}<0$, and elliptic-hyperbolic if and only if $\lambda_{2}>0$. Finally, we notice that $\mathbf{D}$ is the Hessian matrix of $H|_{J=0}$, so $(0,0)$ is an elliptic (hyperbolic) singular point in the Morse sense if and only if $\det(\mathbf{D})=-\lambda_{2}>0$ ($\det(\mathbf{D})=-\lambda_{2}<0$). The claim follows immediately. ∎
Recall that in a semitoric transition family, the type of singularity of the transition point changes as the parameter varies. The following proposition gives a criterion for when a transition point exists and proves that the transition times are roots of a quadratic polynomial.
@@
Let $J=\frac{1}{2}|z_{1}|^{2}-\frac{1}{2}|z_{2}|^{2}$. Fix $\mu_{1},\mu_{2},\mu_{3},\psi,\nu_{2},\nu_{3}\in{\mathbb{R}}$ and $Q_{3},R_{3}\in\mathcal{C}^{\infty}({\mathbb{C}}^{2},{\mathbb{R}})$ with $Q_{3}(z_{1},z_{2})=O(3),R_{3}(z_{1},z_{2})=O(3)$ and $\{J,Q_{3}\}=\{J,R_{3}\}=0$. Define $H_{0}(z_{1},z_{2}):=\nu_{2}|z_{1}|^{2}+\nu_{3}|z_{2}|^{2}+Q_{3}(z_{1},z_{2})$ and $H_{1}(z_{1},z_{2}):=\mu_{1}\mathfrak{R}(e^{i\psi}z_{1}z_{2})+\mu_{2}|z_{1}|^{2}+\mu_{3}|z_{2}|^{2}+R_{3}(z_{1},z_{2})$. Suppose the following hold:
@@
  • $H_{1}(0,0)$ is a non-degenerate local maximum in the Morse-Bott sense;
@@
  • of elliptic-elliptic type when $t\in[0,t^{-})\cup(t^{+},1]$,
  • of focus-focus type when $t\in(t^{-},t^{+})$, and
@@
Since $H_{0}$ has a non-degenerate local minimum (in the Morse-Bott sense) at the origin, the quadratic form $\nu_{2}|z_{1}|^{2}+\nu_{3}|z_{2}|^{2}$ must be positive semi-definite and nonzero. This implies that $\nu_{2}+\nu_{3}>0$. Moreover, since $H_{1}$ has a non-degenerate local maximum (in the Morse-Bott sense) at the origin, $\operatorname{Hess}(H_{1})$ has non-positive eigenvalues with at least one strictly negative eigenvalues, so the sum of eigenvalues of $\operatorname{Hess}(H_{1})$ is strictly negative. Since $\text{det}\big(\lambda I-\text{Hess}(H_{1})\big)=\big(\lambda^{2}-2(\mu_{2}+\mu_{3})\lambda+4\mu_{2}\mu_{3}-\mu_{1}^{2}\big)^{2}$, it follows that $\mu_{2}+\mu_{3}<0$.
Consider the function $H_{t}=(1-t)H_{0}+tH_{1}=t\mu_{1}\mathfrak{R}(e^{i\psi}z_{1}z_{2})+\big(t\mu_{2}+(1-t)\nu_{2}\big)|z_{1}|^{2}+\big(t\mu_{3}+(1-t)\nu_{3}\big)|z_{2}|^{2}+O(3)$. When $t\neq 0$, we have $t\mu_{1}\neq 0$, so by Lemma 3.4 $(J,H_{t})$ is integrable for $t\in(0,1]$. Since $\nu_{2}+\nu_{3}>0$, $(J,H_{0})$ is integrable by the same lemma. Moreover, by Lemma 3.4 the singularity type of $(J,H_{t})$ at $(0,0)$ is determined entirely by the sign of
@@
Indeed, the point is elliptic–elliptic whenever $f(t)>0$, focus–focus whenever $f(t)<0$, and degenerate whenever $f(t)=0$. Thus, the problem is reduced to analyzing the function $f(t)$ on the interval $[0,1]$.
@@
By (1),(3), and (4), the coefficient of $t^{2}$ in $f(t)$, which is $\big((\mu_{2}+\mu_{3})-(\nu_{2}+\nu_{3})\big)^{2}-\mu_{1}^{2}=\{(\mu_{2}+\mu_{3})^{2}-\mu_{1}^{2}\}-2(\mu_{2}+\mu_{3})(\nu_{2}+\nu_{3})+(\nu_{2}+\nu_{3})^{2}$, is strictly positive. By (2) and (3), its discriminant $\Delta=4\mu_{1}^{2}(\nu_{2}+\nu_{3})^{2}$ is strictly positive. Hence $f(t)$ is a concave up parabola with two distinct real roots. Finally, we check their position relative to the interval $[0,1]$. At the left endpoint, by (3) and (4), we have
@@
At the right endpoint, by (1), (3) and (4), we have
@@
Hence, there exist $t^{-},t^{+}$ such that $0<t^{-}<t^{+}<1$ with $f(t^{\pm})=0$, and we have
@@
By Lemma 3.4, $(0,0)$ is a singular point of the elliptic–elliptic type for $t\in[0,t^{-})\cup(t^{+},1]$, focus-focus type for $t\in(t^{-},t^{+})$, and degenerate at $t^{\pm}$. This shows the claimed transition between elliptic-elliptic and focus-focus singularity. ∎

### v2

@@ 2. Polygon spaces
Let $\widetilde{M}:=S^{2}_{r_{1}}\times\cdots\times S^{2}_{r_{n}}$ be the product of symplectic $2$-spheres of radii $r_{1},\ldots,r_{n}$ and let $\mu:\widetilde{M}\to{\mathbb{R}}^{3}$ be the moment map of the diagonal Hamiltonian $SO(3)$-action on $\widetilde{M}$. Under this setting, $\mathcal{P}(r_{1},\ldots,r_{n})=\widetilde{M}\sslash_{0}SO(3)=\mu^{-1}(0)/SO(3)$. Notice that there is a Hamiltonian $S^{1}$-action on $\widetilde{M}$ with moment map $\widetilde{\ell_{12}}(\rho)=|\rho_{1}+\rho_{2}|$, and this $S^{1}$-action commutes with the $SO(3)$-action on $\widetilde{M}$. Hence, we can identify $\ell_{12}^{-1}(c)/S^{1}$ with $F^{-1}(c,0)/(S^{1}\times SO(3))$, where $F=(\widetilde{\ell_{12}},\mu):\widetilde{M}\to{\mathbb{R}}\times{\mathbb{R}}^{3}$. It follows that we can do symplectic reductions by stages: first by the $S^{1}$-action and then by the $SO(3)$-action. In other words, $\ell_{12}^{-1}(c)/S^{1}=(\widetilde{M}\sslash_{c}S^{1})\sslash_{0}SO(3)$.
@@
Let $I,J$ be nonempty subsets of $\{1,\ldots,n\}$. If one of the following holds:
@@
Next, we establish a general criterion characterizing the singular points of $H_{t}$ that will be used in our semitoric families.
@@
Fix $p\in\mathcal{P}(r_{1},\ldots,r_{n})$ and choose a representative $\rho\in\widetilde{\mathcal{P}}(r_{1},\ldots,r_{n})$ of $p$ with $\rho_{1}=(r_{1},0,0)$. For any $t\in[0,1]$, $p$ is a singular point of $H_{t}:=t\ell_{n-2,n-1}^{2}+(1-t)\ell^{2}_{n-1,n}$ if and only if there exists $a\in{\mathbb{R}}$ such that
  • $(a,0,0)\times\rho_{i}=0$ for $i\neq n-2,n-1,n$;
  • $(a,0,0)\times\rho_{n-2}=t\rho_{n-1}\times\rho_{n-2}$; and
@@
Let $r_{1},\dots,r_{5}$ be positive numbers such that the polygon space $M:=\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ is a smooth manifold, and let $J=\ell_{12}:M\to\mathbb{R}$. Let $P,Q$ be pentagons defined in section 1. Assume that the side lengths satisfy the inequalities of one of the cases $(P)$, $(Q)$ or $(P+Q)$ in Theorem 1.1. Then $J$ generates a semi-free Hamiltonian $S^{1}$-action on $M$, and the fixed-point set of this action always contains exactly two connected fixed surfaces $J^{-1}(|r_{1}-r_{2}|)$ and $J^{-1}(r_{1}+r_{2})$.
In addition, depending on which inequalities hold, the action also has isolated fixed points.
  • In case $(P)$, the set of isolated fixed points is $\{P\}$.
  • In case $(Q)$, the set of isolated fixed points is $\{Q\}$.
  • In case $(P+Q)$, the set of isolated fixed points is $\{P,Q\}$.
$J_{\min}$$j_{P}$$J_{\max}$$r_{5}-r_{4}$$r_{3}-J_{\min}$$r_{4}+r_{5}$$J_{\min}$$j_{P}$$J_{\max}$$r_{3}-r_{4}$$r_{5}+J_{\min}$$r_{3}+r_{4}$
(a) Case (P)
$J_{\min}$$j_{Q}$$J_{\max}$$r_{5}-r_{4}$$r_{3}+J_{\min}$$r_{4}+r_{5}$$J_{\min}$$j_{Q}$$J_{\max}$$r_{3}-r_{4}$$r_{5}-J_{\min}$$r_{3}+r_{4}$
(b) Case (Q)
$J_{\min}$$j_{P}$$j_{Q}$$J_{\max}$$r_{5}-r_{4}$$r_{3}-J_{\min}$$r_{3}+J_{\min}$$r_{4}+r_{5}$$J_{\min}$$j_{Q}$$j_{P}$$J_{\max}$$r_{3}-r_{4}$$r_{5}-J_{\min}$$r_{5}+J_{\min}$$r_{3}+r_{4}$
(c) Case (P+Q), where $r_{3}>r_{5}$
Figure 3. Each subfigure contains two moment images of $\mathcal{P}(r_{1},\ldots,r_{5})$ with moment map $(\ell_{12},\ell_{45})$ on the left and moment map $(\ell_{12},\ell_{34})$ on the right, where $r_{1},\ldots,r_{5}$ satisfy inequalities in the corresponding case. In all figures, $J_{\min}=|r_{1}-r_{2}|$, $J_{\max}=r_{1}+r_{2},\ j_{P}=r_{3}+r_{4}-r_{5},\ j_{Q}=-r_{3}+r_{4}+r_{5}$.
@@
The proof of all three cases are nearly identical, so we will only prove the first case. Since in all cases, the inequalities imply $r_{1}\neq r_{2},r_{3}\neq r_{4},r_{4}\neq r_{5}$, so by Lemma 2.3 and [KM96, Proposition 3.11], $(\ell_{12},\ell_{45})$ or $(\ell_{12},\ell_{34})$ is the moment map of an effective Hamiltonian $T^{2}$-action on $M$. It is straightforward to check that the Delzant polygons corresponding to these toric actions are given by Figure 3(a). Since $(\ell_{12},\ell_{34})(P)=(j_{P},r_{3}+r_{4})$, the claims about fixed components follow directly from Delzant’s classification [DEL88] and the local normal form theorem [GS84]. ∎
@@ 3. Semitoric Systems
In this section, we recall the definition of semitoric systems and the classification of non-degenerate singularities. The non-degeneracy of singularities can be defined in any dimension for any integrable system. Since our main interest lies in dimension 4, we will follow Section 2.3 in [LP23] to give the equivalent definition using Hessian matrices. For the definition in any dimension using Cartan subalgebras, we refer the readers to [BF04, Definition 1.23].
@@ 3.1. Singularities of $4$-dimensional integrable systems
  • hyperbolic-hyperbolic if $\lambda_{1},\lambda_{2}>0$; and
@@
Let $p\in M$ be a rank one singular point. In this case, $dJ_{p}$ and $dH_{p}$ are linearly dependent, so the Hamiltonian vector fields $X_{J}(p),X_{H}(p)$ are linearly dependent as well because of the non-degeneracy of $\omega$. Let $L$ be the real span of $X_{J}(p),X_{H}(p)$ and $L^{\omega}$ be the symplectic orthogonal of $L$. Then $A_{\nu,\mu}:=\Omega_{p}^{-1}(\nu\operatorname{Hess}(J)_{p}+\mu\operatorname{Hess}(H)_{p})$ descends to an operator, by slight abuse of notation, still called $A_{\nu,\mu}$ on the 2-dimensional quotient $L^{\omega}/L$, and the eigenvalues of $A_{\nu,\mu}$ are of the form $\pm\lambda$ for $\lambda\in{\mathbb{C}}$ (again, see Proposition 1.2 in [BF04]).
@@
    $J$ is the moment map for an effective Hamiltonian $S^{1}$-action; and
@@ 3.2. Local normal forms of singularities of a semitoric system
In this section, we recall the local theory of singularities of a semitoric system. The best-known local theory is proposed by Eliasson [ELI84]: it is a symplectic Morse lemma for any integrable system. However, since we are studying semitoric systems only, the local normal form given by the Hamiltonian $S^{1}$-action comes in handy, so we will use the following lemma. We refer interested readers to Section 2.2 in [LP23] for a discussion of Eliasson’s normal form and the references therein.
@@ 3.2.1. Rank zero points
[LP23, Proposition 7.5] Let $J=\epsilon(\frac{1}{2}|z_{1}|^{2}-\frac{1}{2}|z_{2}|^{2})$ with $\epsilon\in\{-1,1\}$. Then $H\in\mathcal{C}^{\infty}({\mathbb{C}}^{2},{\mathbb{R}})$ is a Hamiltonian such that $\{J,H\}=0$, $H(0)=0$ and $dH(0)=0$ if and only if there exist $\mu_{1},\mu_{2},\mu_{3},\psi\in{\mathbb{R}}$ and $R_{3}\in\mathcal{C}^{\infty}({\mathbb{C}}^{2},{\mathbb{R}})$ with $R_{3}(z_{1},z_{2})=O(3)$ and $\{J,R_{3}\}=0$ such that
@@
  • of elliptic-elliptic type if $|\mu_{2}+\mu_{3}|>|\mu_{1}|$; and
@@
For a point $p\in M$ fixed by the $S^{1}$-action, the isotropy weights are the weight vectors of the isotropy representation of $S^{1}$ on $T_{p}M$. In dimension $4$, there are two (counted with multiplicity) isotropy weights at each fixed point. The above lemma discusses the case when the isotropy weights are $\pm 1$. In fact, this is the only case in which $p$ can be a focus-focus singularity of the semitoric system. Another case of interest is when $p$ lies on a surface fixed by the $S^{1}$-action. Since the action is effective, the isotropy weights must be $0,1$ or 0,10,-1. We prove the following lemma assuming that $p$ has isotropy weights $0,1$. Notice that we could have used the complex coordinates $z_{1},z_{2}$ and put $J$ as $|z_{1}|^{2}$, but this will put $H$ in a complicated form. Since the non-degeneracy and the type of singularity only depend on the second-order terms in the Taylor expansion, we use the real coordinates to simplify the statement and the proof.
@@
In this case, we have the following matrices:
$\Omega=\begin{pmatrix}\mathbf{J}&\mathbf{O}\\ \mathbf{O}&\mathbf{J}\end{pmatrix},\ \operatorname{Hess}(J)=\begin{pmatrix}2\mathbf{I}&\mathbf{O}\\ \mathbf{O}&\mathbf{O}\end{pmatrix},\ \text{and}\ \operatorname{Hess}(H)=\begin{pmatrix}2a_{11}\mathbf{I}&\mathbf{O}\\ \mathbf{O}&\mathbf{D}\end{pmatrix}$,
where $\mathbf{J}=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$, $\mathbf{D}=\begin{pmatrix}2a_{22}&b_{22}\\ b_{22}&2c_{22}\end{pmatrix}$, $\mathbf{O}$ is the zero matrix, $\mathbf{I}$ is the identity matrix, and $\Omega$ is the matrix of the symplectic form in these coordinates. Fix a number $\lambda\neq-a_{11}$ and consider $A_{\lambda,1}=\Omega^{-1}(\lambda\operatorname{Hess}(J)+\operatorname{Hess}(H))$. The characteristic polynomial of $A_{\lambda,1}$ is $x^{2}+(2\lambda+2a_{11})^{2}][x^{2}+4a_{22}c_{22}-{b_{22}}^{2}]$. Hence, $\chi_{\lambda,1}(X)=(X+(2\lambda+2a_{11})^{2})(X+4a_{22}c_{22}-{b_{22}}^{2})$. The two roots are two real numbers $\lambda_{1}=-(2\lambda+2a_{11})^{2}$ and $\lambda_{2}={b_{22}}^{2}-4a_{22}c_{22}$. Since $\lambda\neq-a_{11}$, we have $\lambda_{1}<0$. By Definition 3.1, $(0,0)$ is elliptic-elliptic if and only if $\lambda_{2}<0$, and elliptic-hyperbolic if and only if $\lambda_{2}>0$. Finally, we notice that $\mathbf{D}$ is the Hessian matrix of $H|_{J=0}$, so $(0,0)$ is an elliptic (respectively, hyperbolic) singular point in the Morse sense if and only if $\det(\mathbf{D})=-\lambda_{2}>0$ (respectively, $\det(\mathbf{D})=-\lambda_{2}<0$). The claim follows immediately. ∎
Recall that in a semitoric transition family, the type of singularity of the transition point changes as the parameter varies. The following proposition establishes a criterion for the existence of a transition point, and shows that the transition times are given by the roots of a quadratic polynomial.
@@
Let $J=\frac{1}{2}|z_{1}|^{2}-\frac{1}{2}|z_{2}|^{2}$. Fix $\mu_{1},\mu_{2},\mu_{3},\psi,\nu_{2},\nu_{3}\in{\mathbb{R}}$ and $Q_{3},R_{3}\in\mathcal{C}^{\infty}({\mathbb{C}}^{2},{\mathbb{R}})$ with $Q_{3}(z_{1},z_{2})=O(3),R_{3}(z_{1},z_{2})=O(3)$ and $\{J,Q_{3}\}=\{J,R_{3}\}=0$. Define $H_{0}(z_{1},z_{2}):=\nu_{2}|z_{1}|^{2}+\nu_{3}|z_{2}|^{2}+Q_{3}(z_{1},z_{2})$ and $H_{1}(z_{1},z_{2}):=\mu_{1}\mathfrak{R}(e^{i\psi}z_{1}z_{2})+\mu_{2}|z_{1}|^{2}+\mu_{3}|z_{2}|^{2}+R_{3}(z_{1},z_{2})$. Suppose the following holds:
@@
  • $H_{1}(0,0)$ is a non-degenerate local maximum in the Morse-Bott sense; and
@@
  • of elliptic-elliptic type when $t\in[0,t^{-})\cup(t^{+},1]$;
  • of focus-focus type when $t\in(t^{-},t^{+})$; and
@@
Since $H_{0}$ has a non-degenerate local minimum (in the Morse-Bott sense) at the origin, the quadratic form $\nu_{2}|z_{1}|^{2}+\nu_{3}|z_{2}|^{2}$ must be positive semi-definite and nonzero. This implies that $\nu_{2}+\nu_{3}>0$. Moreover, since $H_{1}$ has a non-degenerate local maximum (in the Morse-Bott sense) at the origin, $\operatorname{Hess}(H_{1})$ has non-positive eigenvalues with at least one strictly negative eigenvalue, so the sum of eigenvalues of $\operatorname{Hess}(H_{1})$ is strictly negative. Since $\text{det}\big(\lambda I-\text{Hess}(H_{1})\big)=\big(\lambda^{2}-2(\mu_{2}+\mu_{3})\lambda+4\mu_{2}\mu_{3}-\mu_{1}^{2}\big)^{2}$, it follows that $\mu_{2}+\mu_{3}<0$.
Consider the function $H_{t}=(1-t)H_{0}+tH_{1}=t\mu_{1}\mathfrak{R}(e^{i\psi}z_{1}z_{2})+\big(t\mu_{2}+(1-t)\nu_{2}\big)|z_{1}|^{2}+\big(t\mu_{3}+(1-t)\nu_{3}\big)|z_{2}|^{2}+O(3)$. When $t\neq 0$, we have $t\mu_{1}\neq 0$, so by Lemma 3.4 $(J,H_{t})$ is integrable for $t\in(0,1]$. Since $\nu_{2}+\nu_{3}>0$, $(J,H_{0})$ is integrable by the same lemma. Moreover, by Lemma 3.4 the singularity type of $(J,H_{t})$ at $(0,0)$ is entirely determined by the sign of
@@
The point is elliptic–elliptic when $f(t)>0$, focus–focus when $f(t)<0$, and degenerate when $f(t)=0$. Thus, the problem is reduced to analyzing the function $f(t)$ on the interval $[0,1]$.
@@
By (1),(3), and (4), the coefficient of $t^{2}$ in $f(t)$, which is $\big((\mu_{2}+\mu_{3})-(\nu_{2}+\nu_{3})\big)^{2}-\mu_{1}^{2}=\{(\mu_{2}+\mu_{3})^{2}-\mu_{1}^{2}\}-2(\mu_{2}+\mu_{3})(\nu_{2}+\nu_{3})+(\nu_{2}+\nu_{3})^{2}$, is strictly positive. By (2) and (3), its discriminant $\Delta=4\mu_{1}^{2}(\nu_{2}+\nu_{3})^{2}$ is strictly positive. Hence $f(t)$ is a concave up parabola with two distinct real roots. Finally, we check their position relative to the interval $[0,1]$. At the left endpoint $t=0$, by (3) and (4), we have
@@
At the right endpoint $t=1$, by (1), (3), and (4), we have
@@
Hence there exist $t^{-},t^{+}$ such that $0<t^{-}<t^{+}<1$ with $f(t^{\pm})=0$, and we have
@@
By Lemma 3.4, $(0,0)$ is a singular point of elliptic-elliptic type for $t\in[0,t^{-})\cup(t^{+},1]$, of focus-focus type for $t\in(t^{-},t^{+})$, and degenerate at $t^{\pm}$. This shows the claimed transition between elliptic-elliptic and focus-focus singularity. ∎